# Adaptive-RAG: Learning to Adapt Retrieval-Augmented Large Language Models through Question Complexity

### 21 Mar 2024 v1

@@ Abstract
Retrieval-Augmented Large Language Models (LLMs), which incorporate the non-parametric knowledge from external knowledge bases into LLMs, have emerged as a promising approach to enhancing response accuracy in several tasks, such as Question-Answering (QA). However, even though there are various approaches dealing with queries of different complexities, they either handle simple queries with unnecessary computational overhead or fail to adequately address complex multi-step queries; yet, not all user requests fall into only one of the simple or complex categories. In this work, we propose a novel adaptive QA framework, that can dynamically select the most suitable strategy for (retrieval-augmented) LLMs from the simplest to the most sophisticated ones based on the query complexity. Also, this selection process is operationalized with a classifier, which is a smaller LM trained to predict the complexity level of incoming queries with automatically collected labels, obtained from actual predicted outcomes of models and inherent inductive biases in datasets. This approach offers a balanced strategy, seamlessly adapting between the iterative and single-step retrieval-augmented LLMs, as well as the no-retrieval methods, in response to a range of query complexities. We validate our model on a set of open-domain QA datasets, covering multiple query complexities, and show that ours enhances the overall efficiency and accuracy of QA systems, compared to relevant baselines including the adaptive retrieval approaches. Code is available at: https://github.com/starsuzi/Adaptive-RAG.
@@ 1 Introduction
Recent Large Language Models (LLMs) Brown et al. (2020); OpenAI (2023); Touvron et al. (2023); Anil et al. (2023) have shown overwhelming performances across diverse tasks, including question-answering (QA) Yang et al. (2018); Kwiatkowski et al. (2019). However, they still generate factually incorrect answers since their knowledge solely relies on their parametric memory Kasai et al. (2022); Mallen et al. (2023). Meanwhile, memorizing all the (ever-changing) world knowledge may not be possible. To address this problem, retrieval-augmented LLMs Borgeaud et al. (2022); Izacard et al. (2023); Shi et al. (2023), which incorporate non-parametric knowledge into LLMs with additional retrieval modules, have gained increasing attention. Specifically, these models access a knowledge base, serving as an extensive repository of information across various subjects and disciplines, to retrieve relevant information to the given input, and then incorporate this retrieved knowledge into LLMs, which enables the models to stay accurate and current with the world knowledge.
@@
A particularly salient application of retrieval-augmented LLMs is in handling QA tasks, whose goal is to provide correct answers in response to user queries, especially those of high complexity. Early work on retrieval-augmented LLMs focuses primarily on single-hop queries Lazaridou et al. (2022); Ram et al. (2023), where answers are typically found within a single document; therefore, this approach involves retrieving a relevant document based on the query and subsequently integrating this information into QA models to formulate a response. However, unlike this single-hop QA, some queries require connecting and aggregating multiple documents, which are, furthermore, often not answerable through a single-step process of retrieval-and-response. An example query is ‘When did the people who captured Malakoff come to the region where Philipsburg is located?’, which requires four reasoning steps to solve. Therefore, to effectively handle such complex queries, recent studies have concentrated largely on multi-step and multi-reasoning QA, which requires iterative accesses to both LLMs and retrievers multiple times Press et al. (2023); Trivedi et al. (2023), at the cost of heavy computational overheads.
@@
Yet, we should rethink that: In a real-world scenario, are all the requests from users complex? Instead, users might often ask simple and straightforward questions, while only occasionally asking complex ones. Specifically, a query such as ‘Paris is the capital of what?’ is likely to be asked more frequently, compared to the aforementioned multi-step query, and this simpler query might also be easily answerable by the LLMs themselves, without accessing external knowledge. In other words, a multi-step QA approach could result in unnecessary computational overheads for simple queries, even though it is vital for complex queries (see Figure 2 (A)). On the other hand, handling complex queries with single-step-retrieval or even non-retrieval strategies would be largely insufficient (Figure 2 (B)). This suggests a need for an adaptive QA system, which can dynamically adjust the operational strategies of retrieval-augmented LLMs based on the query complexity. While some recent approaches are capable of doing this based on the frequency of entities in queries Mallen et al. (2023) or the generated outputs from models for multi-step QA Trivedi et al. (2023), they are suboptimal: former methods are overly simplistic, failing to consider multi-hop queries; meanwhile, later ones are excessively complex, terminating answer solving steps after several rounds of module accesses.
@@
In this work, considering the diverse complexity levels of real-world queries, we argue that previous one-size-fits-all approaches might be inadequate to cover all of these variations. Instead, we propose to select the most suitable strategy from a range of (retrieval-augmented) LLMs, each of which is tailored to the specific complexity of the input query. Notably, a critical step in this process is pre-defining the query complexity, which is instrumental in determining the most fitting model to it. In this work, we operationalize this process with a novel classifier, which is a smaller model trained to predict the complexity level of incoming queries (see Figure 2 (c)). Moreover, we automatically collect its training datasets without human labeling, by leveraging the predicted outcomes (i.e., which models accurately respond to which queries) as well as by capitalizing on the inherent biases in existing datasets (i.e., samples in the datasets are designed for either single-step or multi-step QA scenarios). This proposed approach can offer a robust middle ground among the iterative LLM augmentation methods for complex queries, the single-step methods for simpler queries, and even no-retrieval-augmented methods for the most straightforward queries (answerable by LLMs themselves), thereby significantly enhancing the overall efficiency and accuracy of QA systems, as shown in Figure 1. We refer to our framework as Adaptive Retrieval-Augmented Generation (Adaptive-RAG).
@@
We validate our Adaptive-RAG using benchmark open-domain QA datasets, covering a wide range of query complexity from single-hop Rajpurkar et al. (2016); Joshi et al. (2017); Kwiatkowski et al. (2019) to multi-hop Yang et al. (2018); Ho et al. (2020); Trivedi et al. (2022b) queries. The experimental results show that our model significantly improves the overall accuracy and efficiency of QA, compared to the prior adaptive retrieval strategies, on multiple LLMs, such as GPT-3.5 Brown et al. (2020) and FLAN-T5 series Chung et al. (2022).
@@
  • We pointed out the realistic scenario of queries of varying complexities, and discovered that existing retrieval-augmented generation approaches tend to be overly simple or complex for them.
  • We propose to adapt retrieval-augmented LLMs to the query complexity assessed by the classifier, which can enable the utilization of the most suitable approach tailored to each query.
@@ 2 Related Work
Multi-hop QA is an extension of conventional Open-domain QA, which additionally requires the system to comprehensively gather and contextualize information from multiple documents (often iteratively), to answer more complex queries Trivedi et al. (2022a); Yang et al. (2018). In the realm of multi-hop QA, the approach to iteratively access both LLMs and the retrieval module is generally employed. Specifically, Khattab et al. (2022)Press et al. (2023)Pereira et al. (2023) and Khot et al. (2023) proposed to first decompose the multi-hop queries into simpler single-hop queries, repeatedly access the LLMs and retriever to solve these sub-queries, and merge their solutions to formulate a complete answer. In contrast to this decomposition-based approach, other recent studies, such as Yao et al. (2023) and Trivedi et al. (2023), explored the interleaving of Chain-of-Thought reasoning Wei et al. (2022b) — a method where a logical sequence of thoughts is generated — with document retrieval, repeatedly applying this process until the reasoning chain generates the answer. In addition, Jiang et al. (2023) introduced an approach for repeatedly retrieving new documents if the tokens within generated sentences have low confidence. However, the aforementioned methods overlooked the fact that, in real-world scenarios, queries are of a wide variety of complexities. Therefore, it would be largely inefficient to iteratively access LLMs and retrievers for every query, which might be simple enough to be solved with a single retrieval step or even only with an LLM itself.
@@
To handle queries of varying complexities, the adaptive retrieval strategy aims to dynamically decide whether to retrieve documents or not, based on each query’s complexity. In this vein, Mallen et al. (2023) proposed to decide the query’s complexity level based on the frequency of its entities and suggested using the retrieval modules only when the frequency falls below a certain threshold. However, this approach, focusing solely on the binary decision to retrieve or not, may not be sufficient for solving more complex queries that require multiple reasoning steps. Additionally, Qi et al. (2021) proposed an approach that performs a fixed set of operations (retrieving, reading, and reranking) multiple times until the answer is derived from the given query, which is built upon traditional BERT-like LMs. However, unlike our Adaptive-RAG which pre-determines the query complexity and adapts the operational behavior of any off-the-shelf LLMs accordingly, this previous approach applies the same fixed operations to every query regardless of its complexity but also necessitates additional specific training to LMs. Concurrent to our work, Asai et al. (2024) suggested training a sophisticated model to dynamically retrieve, critique, and generate the text. Nevertheless, we argue that all the aforementioned adaptive retrieval methods that rely on a single model might be suboptimal in handling a variety of queries of a range of different complexities since they tend to be either overly simple or complex for all the input queries, which gives rise to the new approach that can select the most suitable strategy of retrieval-augmented LLMs tailored to the query complexity.
@@ 3.1 Preliminaries
After the retrieval step is done, we now have a pair of query $\bm{q}$} and its relevant documents $\bm{d}$}. Then, in order to augment LLMs with this retrieved external knowledge, we can incorporate it into the input of LLMs, represented as follows: $\bar{{\bm{a}}}=\texttt{LLM}({\bm{q}},{\bm{d}})$. This process allows LLMs to gain access to external information contained in $\bm{d}$}, which can provide the supplementary context that the internal knowledge of LLM lacks, which can subsequently improve the accurateness and concurrency of LLMs for QA.
@@
In this multi-step approach, LLM interacts with Retriever in several rounds, progressively refining its understanding of $\bm{q}$}, until it formulates the final answer from findings accumulated across those multiple steps. Specifically, the process begins with the initial query $\bm{q}$}, and at every retrieval step $i$, new documents ${\bm{d}}_{i}$ are retrieved from $\mathcal{D}$ and then incorporated into the input of LLMs, as follows: $\bar{{\bm{a}}}_{i}=\texttt{LLM}({\bm{q}},{\bm{d}}_{i},{\bm{c}}_{i})$, where the additional context ${\bm{c}}_{i}$ can be composed of previous documents and outcomes $({\bm{d}}_{1},{\bm{d}}_{2},...,{\bm{d}}_{i-1},\bar{{\bm{a}}}_{1},\bar{{\bm{a}}}_{2},...,\bar{{\bm{a}}}_{i-1})$, and ${\bm{d}}_{i}=\texttt{Retriever}({\bm{q}},{\bm{c}}_{i};D)$ [^1] It is worth noting that implementations of the LLM and retriever vary across different multi-step retrieval-augmented LLM approaches Trivedi et al. (2023); Press et al. (2023); Yao et al. (2023); therefore, the context ${\bm{c}}_{i}$ may incorporate none, some, or all of previous documents and answers.. We would like to note that this iterative, multi-step process enables LLM to construct a more comprehensive and extensive foundation to solve queries effectively, specifically adept at complex multi-hop queries where answers depend on interconnected pieces of information. However, it is important to recognize that this multi-step approach can be resource-intensive due to the repeated accesses to Retriever and LLM, which entail substantial computational costs.
@@ 3.2 Adaptive-RAG: Adaptive Retrieval-Augmented Generation
Note that in real-world scenarios, not all $\bm{q}$} from users have the same level of complexity, necessitating tailored strategies for handling each query. In other words, employing the most basic, non-retrieval-based approach $\texttt{LLM}({\bm{q}})$ to respond to the complex query $\bm{q}$} would be also ineffective (Figure 2, A); conversely, using a more elaborate, multi-step approach $\texttt{LLM}({\bm{q}},{\bm{d}},{\bm{c}})$ for simple $\bm{q}$} would be inefficient (Figure 2, B). Therefore, our adaptive framework is designed to dynamically adjust the query-handling strategy of retrieval-augmented LLMs, which is achieved by pre-determining the complexity of each query prior to attempting a solution. Notably, this framework can offer a robust middle ground with a range of solutions, from the simplest approach for the most straightforward queries, to the one-step approach for moderate queries, and up to the most comprehensive and rigorous approach for complex queries. In addition, since the operations of LLM and Retriever remain consistent regardless of inputs to them, our method can seeminglessly go back and forth across queries of different complexities, without changing the internal model architectures or parameters during adaption.
@@
However, this labeling strategy has a limitation in that not all the queries are assigned the labels, since all three retrieval-augmented approaches may fail to generate the correct answer. On the other hand, the benchmark datasets may already have meaningful inductive biases about the most appropriate retrieval-augmented LLM strategies for their queries, considering the ways they are created (e.g., QA datasets that require sequential reasoning usually necessitate the multi-step approach; while queries of those with the labeled single documents can be ideally answerable with the single-step approach). Therefore, for those queries that remain unlabeled after the first labeling step, we assign ‘B’ to queries in single-hop datasets and ‘C’ to queries in multi-hop datasets. Finally, we train Classifier with these automatically-collected query-complexity pairs [^3] As we automatically assign classifier labels, there might be errors in labeling and might be more advanced strategies to automatically assign labels, which we leave as future work., by using a cross-entropy loss. Then, at inference, we can determine the complexity of the query, which is one of {‘A’, ‘B’, ‘C’}, by forwarding it to Classifier: $o=\texttt{Classifier}({\bm{q}})$.

### 28 Mar 2024 v2

@@ Abstract
Retrieval-Augmented Large Language Models (LLMs), which incorporate the non-parametric knowledge from external knowledge bases into LLMs, have emerged as a promising approach to enhancing response accuracy in several tasks, such as Question-Answering (QA). However, even though there are various approaches dealing with queries of different complexities, they either handle simple queries with unnecessary computational overhead or fail to adequately address complex multi-step queries; yet, not all user requests fall into only one of the simple or complex categories. In this work, we propose a novel adaptive QA framework that can dynamically select the most suitable strategy for (retrieval-augmented) LLMs from the simplest to the most sophisticated ones based on the query complexity. Also, this selection process is operationalized with a classifier, which is a smaller LM trained to predict the complexity level of incoming queries with automatically collected labels, obtained from actual predicted outcomes of models and inherent inductive biases in datasets. This approach offers a balanced strategy, seamlessly adapting between the iterative and single-step retrieval-augmented LLMs, as well as the no-retrieval methods, in response to a range of query complexities. We validate our model on a set of open-domain QA datasets, covering multiple query complexities, and show that ours enhances the overall efficiency and accuracy of QA systems, compared to relevant baselines including the adaptive retrieval approaches. Code is available at: https://github.com/starsuzi/Adaptive-RAG.
@@ 1 Introduction
Recent Large Language Models (LLMs) Brown et al. (2020); OpenAI (2023); Touvron et al. (2023); Anil et al. (2023) have shown overwhelming performances across diverse tasks, including question-answering (QA) Yang et al. (2018); Kwiatkowski et al. (2019). However, they still generate factually incorrect answers since their knowledge solely relies on their parametric memory Kasai et al. (2022); Mallen et al. (2023). Meanwhile, memorizing all the (ever-changing) world knowledge may not be possible. To address this problem, retrieval-augmented LLMs Borgeaud et al. (2022); Izacard et al. (2023); Shi et al. (2023), which incorporate non-parametric knowledge into LLMs with additional retrieval modules, have gained much increasing attention. Specifically, these models access a knowledge base, which serves as an extensive repository of information across various subjects and disciplines, to retrieve information relevant to the given input, and then incorporate the retrieved information into LLMs, which enables them to stay accurate and current with the world knowledge.
@@
A particularly salient application of retrieval-augmented LLMs is to handling QA tasks, whose goal is to provide correct answers in response to user queries, especially those of high complexity. Early work on retrieval-augmented LLMs focuses primarily on single-hop queries Lazaridou et al. (2022); Ram et al. (2023), whose answers are typically found within a single document; therefore, this approach involves retrieving a relevant document based on the query and subsequently integrating this information into QA models to formulate a response. However, unlike this single-hop QA, some queries require connecting and aggregating multiple documents, which are, furthermore, often not answerable through a single-step process of retrieval-and-response. An example query is ‘When did the people who captured Malakoff come to the region where Philipsburg is located?’, which requires four reasoning steps to solve. Therefore, to effectively handle such complex queries, recent studies have concentrated largely on multi-step and multi-reasoning QA, which requires iterative accesses to both LLMs and retrievers multiple times Press et al. (2023); Trivedi et al. (2023), at the cost of heavy computational overheads.
@@
Yet, we should rethink: In a real-world scenario, are all the requests from users complex? Instead, users might often ask simple and straightforward questions, while only occasionally asking complex ones. Specifically, a query such as ‘Paris is the capital of what?’ is likely to be asked more frequently, compared to the aforementioned multi-step query, and this simpler query might also be easily answered by the LLMs themselves, without accessing external knowledge. In other words, a multi-step QA approach could give rise to unnecessary computational overhead for simple queries, even though it would be vital for complex queries (see Figure 2 (A)). On the other hand, handling complex queries with single-step-retrieval or even non-retrieval strategies would be largely insufficient (Figure 2 (B)). This suggests the need for an adaptive QA system, which can dynamically adjust the operational strategies of retrieval-augmented LLMs based on the query complexity. While some recent approaches are capable of doing this based on the frequency of entities in queries Mallen et al. (2023) or on the generated outputs from models for multi-step QA Trivedi et al. (2023), they are still suboptimal: the former methods are overly simplistic, failing to consider multi-hop queries; meanwhile, the latter are excessively complex, terminating answer solving steps after several rounds of module access.
@@
In this work, considering diverse complexity levels of real-world queries, we argue that previous one-size-fits-all approaches might be inadequate to cover all of them. Instead, we propose to select the most suitable strategy from a range of (retrieval-augmented) LLMs, each of which is tailored to the specific complexity of the input query. Notably, a critical step in this process is pre-defining the query complexity, which is instrumental in determining the most fitting model to it. In this work, we operationalize this process with a novel classifier, which is a smaller model trained to predict the complexity level of incoming queries (see Figure 2 (c)). Moreover, we automatically collect its training datasets without human labeling, by leveraging the predicted outcomes (i.e., which models accurately respond to which queries) as well as by capitalizing on the inherent biases in existing datasets (i.e., samples in the datasets are designed either for single-step or for multi-step QA scenarios). This proposed method can offer a robust middle ground among the iterative LLM augmentation methods for complex queries, single-step methods for simpler queries, and even no-retrieval-augmented methods for the most straightforward queries (answerable by LLMs themselves), thus significantly enhancing the overall efficiency and accuracy, as shown in Figure 1. We refer to our framework as Adaptive Retrieval-Augmented Generation (Adaptive-RAG).
@@
We validate Adaptive-RAG using benchmark open-domain QA datasets, covering a wide range of query complexity from single-hop Rajpurkar et al. (2016); Joshi et al. (2017); Kwiatkowski et al. (2019) to multi-hop Yang et al. (2018); Ho et al. (2020); Trivedi et al. (2022b) queries. The experimental results show that ours significantly improves the overall accuracy and efficiency, compared to the prior adaptive strategies, on multiple LLMs, such as GPT-3.5 Brown et al. (2020) and FLAN-T5 series Chung et al. (2022).
@@
  • We point out the realistic scenario of queries of varying complexities, and find out that existing retrieval-augmented generation approaches tend to be overly simple or complex.
  • We adapt retrieval-augmented LLMs to the query complexity assessed by the classifier, which enables the utilization of the most suitable approach tailored to each query.
@@ 2 Related Work
Multi-hop QA is an extension of conventional Open-domain QA, which additionally requires the system to comprehensively gather and contextualize information from multiple documents (often iteratively), to answer more complex queries Trivedi et al. (2022a); Yang et al. (2018). In the realm of multi-hop QA, the approach to iteratively access both LLMs and the retrieval module is generally employed. Specifically, Khattab et al. (2022)Press et al. (2023)Pereira et al. (2023) and Khot et al. (2023) proposed to first decompose the multi-hop queries into simpler single-hop queries, repeatedly access the LLMs and retriever to solve these sub-queries, and merge their solutions to formulate a complete answer. In contrast to this decomposition-based approach, other recent studies, such as Yao et al. (2023) and Trivedi et al. (2023), explored the interleaving of Chain-of-Thought reasoning Wei et al. (2022b) — a method where a logical sequence of thoughts is generated — with document retrieval, repeatedly applying this process until the reasoning chain generates the answer. In addition, Jiang et al. (2023) introduced an approach to repeatedly retrieving new documents if the tokens within generated sentences have low confidence. However, the aforementioned methods overlooked the fact that, in real-world scenarios, queries are of a wide variety of complexities. Therefore, it would be largely inefficient to iteratively access LLMs and retrievers for every query, which might be simple enough with a single retrieval step or even only with an LLM itself.
@@
To handle queries of varying complexities, the adaptive retrieval strategy aims to dynamically decide whether to retrieve documents or not, based on each query’s complexity. In this vein, Mallen et al. (2023) proposed to decide the query’s complexity level based on the frequency of its entities and suggested using the retrieval modules only when the frequency falls below a certain threshold. However, this approach, focusing solely on the binary decision of whether to retrieve or not, may not be sufficient for more complex queries that require multiple reasoning steps. Additionally, Qi et al. (2021) proposed an approach that performs a fixed set of operations (retrieving, reading, and reranking) multiple times until the answer is derived for the given query, which is built upon traditional BERT-like LMs. However, unlike our Adaptive-RAG which pre-determines the query complexity and adapts the operational behavior of any off-the-shelf LLMs accordingly, this approach applies the same fixed operations to every query regardless of its complexity but also necessitates additional specific training to LMs. Concurrent to our work, Asai et al. (2024) suggested training a sophisticated model to dynamically retrieve, critique, and generate the text. Nevertheless, we argue that all the aforementioned adaptive retrieval methods that rely on a single model might be suboptimal in handling a variety of queries of a range of different complexities since they tend to be either overly simple or complex for all the input queries, which demands a new approach that can select the most suitable strategy of retrieval-augmented LLMs tailored to the query complexity.
@@ 3.1 Preliminaries
After the retrieval step is done, we now have a pair of query $\bm{q}$} and its relevant documents $\bm{d}$}. Then, in order to augment LLMs with this retrieved external knowledge, we can incorporate it into the input of LLMs, represented as follows: $\bar{{\bm{a}}}=\texttt{LLM}({\bm{q}},{\bm{d}})$. This process allows LLMs to gain access to external information contained in $\bm{d}$}, which can provide the supplementary context that the internal knowledge of LLM lacks, which can subsequently improve the accuracy and concurrency of LLMs for QA.
@@
In this multi-step approach, LLM interacts with Retriever in several rounds, progressively refining its understanding of $\bm{q}$}, until it formulates the final answer from findings accumulated across these multiple steps. Specifically, the process begins with the initial query $\bm{q}$}, and at every retrieval step $i$, new documents ${\bm{d}}_{i}$ are retrieved from $\mathcal{D}$ and then incorporated into the input of LLMs, as follows: $\bar{{\bm{a}}}_{i}=\texttt{LLM}({\bm{q}},{\bm{d}}_{i},{\bm{c}}_{i})$, where the additional context ${\bm{c}}_{i}$ can be composed of previous documents and outcomes $({\bm{d}}_{1},{\bm{d}}_{2},...,{\bm{d}}_{i-1},\bar{{\bm{a}}}_{1},\bar{{\bm{a}}}_{2},...,\bar{{\bm{a}}}_{i-1})$, and ${\bm{d}}_{i}=\texttt{Retriever}({\bm{q}},{\bm{c}}_{i};D)$ [^1] It is worth noting that implementations of the LLM and retriever vary across different multi-step retrieval-augmented LLM approaches Trivedi et al. (2023); Press et al. (2023); Yao et al. (2023); therefore, the context ${\bm{c}}_{i}$ may incorporate none, some, or all of the previous documents and answers.. We would like to note that this iterative, multi-step process enables LLM to construct a more comprehensive and extensive foundation to solve queries effectively, specifically adept at complex multi-hop queries where answers depend on interconnected pieces of information. However, it is important to recognize that this multi-step approach can be resource-intensive due to the repeated accesses to Retriever and LLM, which entail substantial computational costs.
@@ 3.2 Adaptive-RAG: Adaptive Retrieval-Augmented Generation
Note that in real-world scenarios, not all $\bm{q}$} from users have the same level of complexity, necessitating tailored strategies for handling each query. In other words, employing the most basic, non-retrieval-based approach $\texttt{LLM}({\bm{q}})$ to respond to the complex query $\bm{q}$} would be also ineffective (Figure 2, A); conversely, using a more elaborate multi-step approach $\texttt{LLM}({\bm{q}},{\bm{d}},{\bm{c}})$ for simple $\bm{q}$} would be inefficient (Figure 2, B). Therefore, our adaptive framework is designed to dynamically adjust the query-handling strategy of retrieval-augmented LLMs, which is achieved by determining the complexity of each query before attempting a solution. Notably, this framework can offer a robust middle ground with a range of solutions, from the simplest approach for the most straightforward queries, to the one-step approach for moderate queries, and up to the most comprehensive and rigorous approach for complex queries. In addition, since the operations of LLM and Retriever remain consistent regardless of inputs to them, our method can seeminglessly go back and forth across queries of different complexities, without changing the internal model architecture or parameters during adaption.
@@
However, this labeling strategy has a limitation in that not all the queries are assigned labels, since the three retrieval-augmented approaches may all fail to generate the correct answer. On the other hand, the benchmark datasets may already have meaningful inductive biases about the most appropriate retrieval-augmented LLM strategies for their queries, considering the ways they are created (e.g., QA datasets that require sequential reasoning usually necessitate a multi-step approach; while queries of those with labeled single documents can be ideally answerable with the single-step approach). Therefore, for those queries that remain unlabeled after the first labeling step, we assign ‘B’ to queries in single-hop datasets and ‘C’ to queries in multi-hop datasets. Finally, we train Classifier with these automatically-collected query-complexity pairs [^3] As we automatically assign classifier labels, there might be errors in labeling and might be more advanced strategies to automatically assign labels, which we leave as future work., by using a cross-entropy loss. Then, at inference, we can determine the complexity of the query, which is one of {‘A’, ‘B’, ‘C’}, by forwarding it to Classifier: $o=\texttt{Classifier}({\bm{q}})$.